# Parametric Algorithms for the 5-Modular Analog of ES (Sierpiński): Structure of Solutions, Parameterization, and Constructive Proofs (SERP)

### v1

@@ Abstract
This study continues the work for coefficient 44 (the Erdős–Straus conjecture) [Dyachenko_ESC_ED2]; here a similar structure of parameterization and solutions is extended to coefficient 55. In the analytic applications, averaging tools are provided, used for density estimates in parametric boxes.
@@ 1 Introduction
Analogous to the Erdős–Straus conjecture for $\tfrac{4}{P}$, here the coefficient 55 leads to a change in the structure of parameterizations and to new types of solutions. For primes $P\equiv 1\pmod{5}$ two constructive classes are distinguished:
@@ 2 Motivation
The possibility of transferring the constructive method from coefficient 44 to coefficient 55 was correctly noted in [Mballa_ESC_SERP]. In the present work this transfer is carried out explicitly within the framework of the $\mathrm{ED1}/\mathrm{ED2}$ methods developed in [Dyachenko_ESC_ED2]. The key element here is the mechanism of finite covering of residue classes, which guarantees the existence of solutions for every prime $P\equiv 1\pmod{5}$ without recourse to factorization. Thus, the motivation of the study is to show that constructive methods based on lattices and covering of residue classes allow one to move from asymptotic reasoning to effective algorithms with polylogarithmic complexity.
@@ 3.1 Explicit decompositions for $P\not\equiv 1\pmod{5}$
$\frac{5}{P}=\frac{1}{\frac{P+4}{5}}+\frac{1}{2(P^{\prime}+1)(5P^{\prime}+4)}+\frac{1}{2(P^{\prime}+1)(5P^{\prime}+4)}.$
@@
$\frac{5}{P}=\frac{1}{\frac{P+4}{5}}+\frac{1}{(P^{\prime}+1)(5P^{\prime}+3)}+\frac{1}{(P^{\prime}+1)(5P^{\prime}+3)}.$
@@
$\frac{5}{P}=\frac{1}{\frac{P+4}{5}}+\frac{1}{(P^{\prime}+1)(5P^{\prime}+4)}+\frac{1}{(P^{\prime}+1)(5P^{\prime}+4)},$
@@ 6.1 Setup and identities
Multiplying by AbcPAbcP, we obtain
@@ 7 Multiplicity configurations with respect to $P$ and classification
Multiplying (1.1) by ABCPABCP: $5ABC=P(AB+AC+BC)$. Modulo $P$: $5ABC\equiv 0$, hence $P\mid ABC$. ∎
@@
The minimal denominator AA is not divisible by $P$.
@@ 7.3 Box threshold via AA and $P$
it follows that the minimal denominator AA always lies in the interval (P/5, 3P/5)(P/5,\;3P/5). This provides natural bounds for the parametric box.
@@ 8.1 Setup and kernel, fully analogous to ESC
Multiplying by AbcPAbcP and introducing the parameter $\delta\in\mathbb{N}$, we obtain
@@ 9 Connection $\mathrm{ED2}\leftrightarrow\mathrm{ED1}$: convolution and anticonvolution
A detailed analysis of the convolution and anticonvolution mechanism for the case $k=4$ is given in the work on the Erdős–Straus conjecture [Dyachenko_ESC_ED2]; here we transfer the same construction to coefficient 55.
@@ Appendix B Local applications
By the second condition $r\mid(5P\delta+1)$, hence $s\in\mathbb{N}$. Modulo 55: $5P\delta+1\equiv 1$, and $r\equiv 4$ is invertible, with $r^{-1}\equiv 4\pmod{5}$. Thus $s\equiv 1\cdot r^{-1}\equiv 4\pmod{5}$. ∎

### v2

@@ Abstract
This study continues the work for coefficient $4$ (the Erdős–Straus conjecture) [Dyachenko_ESC_ED2]; here a similar structure of parameterization and solutions is extended to coefficient $5$. In the analytic applications, averaging tools are provided, used for density estimates in parametric boxes.
@@ 1 Introduction
Analogous to the Erdős–Straus conjecture for $\tfrac{4}{P}$, here the coefficient $5$ leads to a change in the structure of parameterizations and to new types of solutions. For primes $P\equiv 1\pmod{5}$ two constructive classes are distinguished:
@@ 2 Motivation
The possibility of transferring the constructive method from coefficient $4$ to coefficient $5$ was correctly noted in [Mballa_ESC_SERP]. In the present work this transfer is carried out explicitly within the framework of the $\mathrm{ED1}/\mathrm{ED2}$ methods developed in [Dyachenko_ESC_ED2]. The key element here is the mechanism of finite covering of residue classes, which guarantees the existence of solutions for every prime $P\equiv 1\pmod{5}$ without recourse to factorization. Thus, the motivation of the study is to show that constructive methods based on lattices and covering of residue classes allow one to move from asymptotic reasoning to effective algorithms with polylogarithmic complexity.
@@ 3.1 Explicit decompositions for $P\not\equiv 1\pmod{5}$
$\frac{5}{P}=\frac{5}{5P^{\prime}+4}=\frac{1}{P^{\prime}+1}+\frac{1}{2\cdot(P^{\prime}+1)\cdot(5P^{\prime}+4)}+\frac{1}{2\cdot(P^{\prime}+1)\cdot(5P^{\prime}+4)}.$
@@
$\frac{5}{P}=\frac{1}{P^{\prime}+1}+\frac{1}{(P^{\prime}+1)(5P^{\prime}+3)}+\frac{1}{(P^{\prime}+1)(5P^{\prime}+3)}.$
@@
$\frac{5}{P}=\frac{5}{5P^{\prime}+2}=\frac{1}{P^{\prime}+1}+\frac{2}{(P^{\prime}+1)\cdot(5P^{\prime}+2)}+\frac{1}{(P^{\prime}+1)\cdot(5P^{\prime}+2)}=\frac{1}{P^{\prime}+1}+\frac{1}{\frac{(P^{\prime}+1)}{2}\cdot(5P^{\prime}+2)}+\frac{1}{(P^{\prime}+1)\cdot(5P^{\prime}+2)},$
@@ 6.1 Setup and identities
Multiplying by $AbcP$, we obtain
@@ 7 Multiplicity configurations with respect to $P$ and classification
Multiplying (1.1) by $ABCP$: $5ABC=P(AB+AC+BC)$. Modulo $P$: $5ABC\equiv 0$, hence $P\mid ABC$. ∎
@@
The minimal denominator $A$ is not divisible by $P$.
@@ 7.3 Box threshold via $A$ and $P$
it follows that the minimal denominator $A$ always lies in the interval (P/5, 3P/5)(P/5,\;3P/5). This provides natural bounds for the parametric box.
@@ 8.1 Setup and kernel, fully analogous to ESC
Multiplying by $AbcP$ and introducing the parameter $\delta\in\mathbb{N}$, we obtain
@@ 9 Connection $\mathrm{ED2}\leftrightarrow\mathrm{ED1}$: convolution and anticonvolution
A detailed analysis of the convolution and anticonvolution mechanism for the case $k=4$ is given in the work on the Erdős–Straus conjecture [Dyachenko_ESC_ED2]; here we transfer the same construction to coefficient $5$.
@@ Appendix B Local applications
By the second condition $r\mid(5P\delta+1)$, hence $s\in\mathbb{N}$. Modulo $5$: $5P\delta+1\equiv 1$, and $r\equiv 4$ is invertible, with $r^{-1}\equiv 4\pmod{5}$. Thus $s\equiv 1\cdot r^{-1}\equiv 4\pmod{5}$. ∎